# TENG++: Time-Evolving Natural Gradient for Solving PDEs With Deep Neural Nets under General Boundary Conditions

### v2

@@ 1 Introduction
Partial Differential Equations (PDEs) are foundational to modeling a wide range of physical, biological, and engineering systems. Solving PDEs is a cornerstone of computational science, but traditional numerical methods such as finite element and finite difference approaches can become computationally prohibitive, particularly in high-dimensional or complex domains. In recent years, Physics-Informed Neural Networks (PINNs)(Raissi:PINN) have emerged as a promising alternative, leveraging the expressive power of deep learning to approximate solutions to PDEs by embedding the governing equations into the network’s loss function.
@@
Physics-Informed Neural Networks (PINNs) offer significant advantages, including being mesh-free and computationally efficient. These qualities make them particularly appealing for solving complex systems of equations like the Navier-Stokes equations, which have been a major focus of research in fluid dynamics. Despite significant progress over the past 50 years in the mathematical community, challenges remain: seamlessly incorporating noisy data into existing algorithms remains elusive, and generating meshes in high-dimensional spaces is inherently complex. In (Cai:mesh_free), the authors leveraged PINNs to address these challenges, demonstrating their effectiveness in solving a range of fluid dynamics problems. Remarkably, PINNs can also tackle ill-posed problems (Cai:ill_posed), which is typically infeasible using classical methods.
@@
Despite their advantages of being mesh-free and computationally efficient, Physics-Informed Neural Networks (PINNs) have yet to achieve the level of accuracy offered by classical methods (Grossmann:accuracy; Cuomo:accuracy). This limitation is particularly evident in solving initial value problems in partial differential equations (PDEs), which are fundamental for modeling the evolution of dynamical systems. While machine learning techniques have shown promise in approximating solutions to complex partial differential equations (PDEs), they often struggle to achieve high accuracy, particularly when dealing with intricate initial conditions. This difficulty arises from the cumulative propagation of errors in PDE solvers over time, necessitating precise solutions at each time step to maintain overall accuracy.
@@
Although various training strategies, such as global-in-time training (Muller2023) and sequential-in-time training (Chen2023a; Berman2023), have been developed to mitigate these challenges, achieving consistent accuracy remains a significant obstacle to the practical application of PINNs. Addressing this, the authors in (Chen:TENG) proposed the Time-Evolving Natural Gradient (TENG) method, which outperformed all previously benchmarked approaches, offering a promising advancement in the field.
@@
In (Chen:TENG), the authors primarily focused on periodic boundary conditions. While effective in certain contexts, periodic boundary conditions are limited in scope. Non-periodic boundary conditions, such as Dirichlet or Neumann conditions, are essential for modeling a broader range of real-world phenomena where solutions are governed by fixed values or fluxes at the boundaries. These conditions enable accurate representation of complex systems, including heat transfer, fluid flow, and electromagnetic fields, where periodic assumptions are not applicable.
@@
In this work, we extend the method proposed in (Chen:TENG) to handle arbitrary boundary conditions, broadening its applicability to a larger class of PDE problems. This enhancement significantly increases its potential for addressing real-world challenges. To validate the effectiveness of our approach, we present a series of numerical experiments on benchmark PDE problems of varying complexity.
@@ 2.1 Machine Learning for Solving PDEs
Global-in-time optimization, such as the Physics-Informed Neural Network (PINN) framework (Raissi2019), aims to optimize a neural network representation of the solution over the entire time and spatial domain simultaneously. Variants of this approach include the Deep Ritz Method (Weinan2017), which leverages the variational form of the PDE when it exists. While effective in many cases, these methods can struggle with maintaining accuracy, particularly in problems with intricate boundary conditions, where incorporating boundary constraints seamlessly remains challenging.
In contrast, sequential-in-time optimization, also known as the neural Galerkin method, focuses on updating the neural network representation step-by-step in time. Methods in this category include the Time-Dependent Variational Principle (TDVP) (Dirac1930; Koch2007; Carleo2017; Du2021; Berman2023) and Optimization-Based Time Integration (OBTI) (Kochkov2018; Gutierrez2022). Sequential-in-time strategies are particularly advantageous in handling evolving boundary conditions, as they allow for explicit incorporation of such conditions at each time step.
Beyond these direct PDE-solving strategies, data-driven approaches have also been employed to model PDEs based on observed data. These methods include Neural ODEs (Chen2018), Graph Neural Networks (Pfaff2020; SanchezGonzalez2020), Neural Fourier Operators (Li2020), and DeepONet (Lu2019). While these techniques have demonstrated strong performance in approximating PDE solutions, adapting them to accurately and efficiently enforce complex boundary conditions remains a significant challenge.
@@ 2.2 Natural Gradient Optimization
Natural gradient optimization, introduced by Amari (Amari1998), has become a cornerstone in second-order optimization techniques for machine learning. By adjusting the gradient direction based on the Fisher information matrix, natural gradient descent incorporates the geometry of the parameter space, leading to faster and more stable convergence compared to standard gradient methods. This feature makes it particularly effective for problems with complex constraints, such as boundary conditions in PDEs.
Natural gradient methods have been widely applied in neural network optimization (Pascanu2013), reinforcement learning (Kakade2001), quantum optimization, and PINN training (Muller2023). In the context of solving PDEs, these methods can improve the enforcement of boundary conditions by effectively navigating the optimization landscape shaped by the PDE constraints and boundary terms. Recent works have demonstrated the utility of natural gradients in PINN training (Muller2023), highlighting their potential to address issues such as cumulative error propagation and the incorporation of boundary constraints.
@@ 3 Methodology
The Time-Evolving Natural Gradient (TENG) framework (Chen:TENG) combines the time-dependent variational principle with optimization-based time integration using the natural gradient. This approach leverages second-order optimization to enhance convergence and numerical stability. While TENG has demonstrated effectiveness under periodic boundary conditions, it does not generalize to problems with arbitrary boundary conditions, such as Dirichlet or Neumann conditions. In this work, we extend the TENG framework to address these limitations.
@@ 3.2 Optimization Process
The details of the parameter update process are implemented in the TENG_stepper subroutine, described in Algorithm 1.
@@
  Compute the gradient: $\Delta u(x)\leftarrow-\alpha_{n}\frac{\partial L(\hat{u}_{\theta},u_{\text{target}})}{\partial\hat{u}_{\theta}}$
  Compute the Jacobian: $J(x)_{j}\leftarrow\frac{\partial\hat{u}_{\theta}(x)}{\partial\theta_{j}}$
@@ 3.3.1 Euler Method
Euler’s method is a first-order integration scheme that updates the solution at each time step based on the PDE residual. The algorithm is summarized in Algorithm 2.
@@
  Compute $u_{\text{target}}(x)\leftarrow\hat{u}_{\theta_{t}}(x)+\Delta tL\hat{u}_{\theta_{t}}(x)$
@@ 3.3.2 Heun Method
Heun’s method is a second-order integration scheme that incorporates an intermediate step to improve accuracy. The algorithm is summarized in Algorithm 3.
@@
  Compute $u_{\text{temp}}(x)\leftarrow\hat{u}_{\theta_{t}}(x)+\Delta tL\hat{u}_{\theta_{t}}(x)$
@@
  Compute $u_{\text{target}}(x)\leftarrow\hat{u}_{\theta_{t}}(x)+\frac{\Delta t}{2}\big(L\hat{u}_{\theta_{t}}(x)+L\hat{u}_{\theta_{\text{temp}}}(x)\big)$
@@ 4.1 Experiment 1
$u_{0}(r,\theta)=\frac{1}{4}(Z_{01}(r,\theta)-\frac{1}{4}Z_{02}(r,\theta)+\frac{1}{16}Z_{03}(r,\theta)\\ -\frac{1}{64}Z_{04}(r,\theta)+Z_{11}(r,\theta)-\frac{1}{2}Z_{12}(r,\theta)+\frac{1}{4}Z_{13}(r,\theta)\\ -\frac{1}{8}Z_{14}(r,\theta)+Z_{21}(r,\theta)+Z_{31}(r,\theta)+Z_{41}(r,\theta))$ (1)
@@ 4.2 Experiment 2
$u_{0}(r,\theta)=\frac{1}{4}(Z_{01}(r,\theta)-\frac{1}{4}Z_{02}(r,\theta)+\frac{1}{16}Z_{03}(r,\theta)\\ -\frac{1}{64}Z_{04}(r,\theta)+Z_{11}(r,\theta)-\frac{1}{2}Z_{12}(r,\theta)+\frac{1}{4}Z_{13}(r,\theta)\\ -\frac{1}{8}Z_{14}(r,\theta)+Z_{21}(r,\theta)+Z_{31}(r,\theta)+Z_{41}(r,\theta))$ (2)
@@ 5 Discussion
In this work, we generalized periodic boundary conditions in (Chen:TENG) to Dirichlet boundary conditions, demonstrating the adaptability of our approach to handle different types of constraints. This extension broadens the applicability of the method to a wider range of physical and mathematical problems, where Dirichlet boundary conditions play a critical role. The results underscore the potential of this generalized framework for accurately solving boundary value problems while maintaining computational efficiency and flexibility.
@@ 6 Acknowledgments
We would like to thank Professor Di Luo and Zhuo Chen, their valuable suggestions and great support throughout this project.

### v3

# Solving PDEs With Deep Neural Nets under General Boundary Conditions

@@ 1 Introduction
Partial Differential Equations (PDEs) are foundational to modeling a wide range of physical, biological, and engineering systems. Solving PDEs is a cornerstone of computational science, but traditional numerical methods such as finite element and finite difference approaches can become computationally prohibitive, particularly in high-dimensional or complex domains. In recent years, Physics-Informed Neural Networks (PINNs)Raissi et al. (2017) have emerged as a promising alternative, leveraging the expressive power of deep learning to approximate solutions to PDEs by embedding the governing equations into the network’s loss function.
@@
Physics-Informed Neural Networks (PINNs) offer significant advantages, including being mesh-free and computationally efficient. These qualities make them particularly appealing for solving complex systems of equations like the Navier-Stokes equations, which have been a major focus of research in fluid dynamics. Despite significant progress over the past 50 years in the mathematical community, challenges remain: seamlessly incorporating noisy data into existing algorithms remains elusive, and generating meshes in high-dimensional spaces is inherently complex. In Cai et al. (2021a), the authors leveraged PINNs to address these challenges, demonstrating their effectiveness in solving a range of fluid dynamics problems. Remarkably, PINNs can also tackle ill-posed problems Cai et al. (2021b), which is typically infeasible using classical methods.
@@
Despite their advantages of being mesh-free and computationally efficient, Physics-Informed Neural Networks (PINNs) have yet to achieve the level of accuracy offered by classical methods Grossmann et al. (2023); Cuomo et al. (2022). This limitation is particularly evident in solving initial value problems in partial differential equations (PDEs), which are fundamental for modeling the evolution of dynamical systems. While machine learning techniques have shown promise in approximating solutions to complex partial differential equations (PDEs), they often struggle to achieve high accuracy, particularly when dealing with intricate initial conditions. This difficulty arises from the cumulative propagation of errors in PDE solvers over time, necessitating precise solutions at each time step to maintain overall accuracy.
@@
Although various training strategies, such as global-in-time training Müller and Zeinhofer (2023) and sequential-in-time training Chen et al. (2023); Berman and Peherstorfer (2023), have been developed to mitigate these challenges, achieving consistent accuracy remains a significant obstacle to the practical application of PINNs. Addressing this, the authors in Chen et al. (2024) proposed the Time-Evolving Natural Gradient (TENG) method, which outperformed all previously benchmarked approaches, offering a promising advancement in the field.
@@
In Chen et al. (2024), the authors primarily focused on periodic boundary conditions. While effective in certain contexts, periodic boundary conditions are limited in scope. Non-periodic boundary conditions, such as Dirichlet or Neumann conditions, are essential for modeling a broader range of real-world phenomena where solutions are governed by fixed values or fluxes at the boundaries. These conditions enable accurate representation of complex systems, including heat transfer, fluid flow, and electromagnetic fields, where periodic assumptions are not applicable.
@@
In this work, we extend the method proposed in Chen et al. (2024) to handle arbitrary boundary conditions, broadening its applicability to a larger class of PDE problems. This enhancement significantly increases its potential for addressing real-world challenges. To validate the effectiveness of our approach, we present a series of numerical experiments on benchmark PDE problems of varying complexity.
@@ 2.1 Machine Learning for Solving PDEs
Global-in-time optimization, such as the Physics-Informed Neural Network (PINN) framework Raissi et al. (2019), aims to optimize a neural network representation of the solution over the entire time and spatial domain simultaneously. Variants of this approach include the Deep Ritz Method E and Yu (2017), which leverages the variational form of the PDE when it exists. While effective in many cases, these methods can struggle with maintaining accuracy, particularly in problems with intricate boundary conditions, where incorporating boundary constraints seamlessly remains challenging.
In contrast, sequential-in-time optimization, also known as the neural Galerkin method, focuses on updating the neural network representation step-by-step in time. Methods in this category include the Time-Dependent Variational Principle (TDVP) Dirac (1930); Koch and Lubich (2007); Carleo and Troyer (2017); Du and Zaki (2021); Berman and Peherstorfer (2023) and Optimization-Based Time Integration (OBTI) Kochkov and Clark (2018); Gutierrez and Mendl (2022). Sequential-in-time strategies are particularly advantageous in handling evolving boundary conditions, as they allow for explicit incorporation of such conditions at each time step.
Beyond these direct PDE-solving strategies, data-driven approaches have also been employed to model PDEs based on observed data. These methods include Neural ODEs Chen et al. (2018), Graph Neural Networks Pfaff et al. (2020); Sanchez-Gonzalez et al. (2020), Neural Fourier Operators Li et al. (2020), and DeepONet Lu et al. (2019). While these techniques have demonstrated strong performance in approximating PDE solutions, adapting them to accurately and efficiently enforce complex boundary conditions remains a significant challenge.
@@ 2.2 Natural Gradient Optimization
Natural gradient optimization, introduced by Amari Amari (1998), has become a cornerstone in second-order optimization techniques for machine learning. By adjusting the gradient direction based on the Fisher information matrix, natural gradient descent incorporates the geometry of the parameter space, leading to faster and more stable convergence compared to standard gradient methods. This feature makes it particularly effective for problems with complex constraints, such as boundary conditions in PDEs.
Natural gradient methods have been widely applied in neural network optimization Pascanu et al. (2013), reinforcement learning Kakade (2001), quantum optimization, and PINN training Müller and Zeinhofer (2023). In the context of solving PDEs, these methods can improve the enforcement of boundary conditions by effectively navigating the optimization landscape shaped by the PDE constraints and boundary terms. Recent works have demonstrated the utility of natural gradients in PINN training Müller and Zeinhofer (2023), highlighting their potential to address issues such as cumulative error propagation and the incorporation of boundary constraints.
@@ 3 Methodology
The Time-Evolving Natural Gradient (TENG) framework Chen et al. (2024) combines the time-dependent variational principle with optimization-based time integration using the natural gradient. This approach leverages second-order optimization to enhance convergence and numerical stability. While TENG has demonstrated effectiveness under periodic boundary conditions, it does not generalize to problems with arbitrary boundary conditions, such as Dirichlet or Neumann conditions. In this work, we extend the TENG framework to address these limitations.
@@ 3.2 Optimization Process
The details of the parameter update process are implemented in the TENG_stepper subroutine, described in 1.
@@
  Compute the gradient: $\Delta u(x)\leftarrow-\alpha_{n}\frac{\partial L(\widehat{u}_{\theta},u_{\text{target}})}{\partial\widehat{u}_{\theta}}$
  Compute the Jacobian: $J(x)_{j}\leftarrow\frac{\partial\widehat{u}_{\theta}(x)}{\partial\theta_{j}}$
@@ 3.3.1 Euler Method
Euler’s method is a first-order integration scheme that updates the solution at each time step based on the PDE residual. The algorithm is summarized in 2.
@@
  Compute $u_{\text{target}}(x)\leftarrow\widehat{u}_{\theta_{t}}(x)+\Delta tL\widehat{u}_{\theta_{t}}(x)$
@@ 3.3.2 Heun Method
Heun’s method is a second-order integration scheme that incorporates an intermediate step to improve accuracy. The algorithm is summarized in 3.
@@
  Compute $u_{\text{temp}}(x)\leftarrow\widehat{u}_{\theta_{t}}(x)+\Delta tL\widehat{u}_{\theta_{t}}(x)$
@@
  Compute $u_{\text{target}}(x)\leftarrow\widehat{u}_{\theta_{t}}(x)+\frac{\Delta t}{2}\big(L\widehat{u}_{\theta_{t}}(x)+L\widehat{u}_{\theta_{\text{temp}}}(x)\big)$
@@ 4.1 Experiment 1
$u_{0}(r,\theta)$ $=\frac{1}{4}(Z_{01}(r,\theta)-\frac{1}{4}Z_{02}(r,\theta)+\frac{1}{16}Z_{03}(r,\theta)$
$-\frac{1}{64}Z_{04}(r,\theta)+Z_{11}(r,\theta)-\frac{1}{2}Z_{12}(r,\theta)+\frac{1}{4}Z_{13}(r,\theta)$
$-\frac{1}{8}Z_{14}(r,\theta)+Z_{21}(r,\theta)+Z_{31}(r,\theta)+Z_{41}(r,\theta))$
@@ 4.2 Experiment 2
$u_{0}(r,\theta)$ $=\frac{1}{4}(Z_{01}(r,\theta)-\frac{1}{4}Z_{02}(r,\theta)+\frac{1}{16}Z_{03}(r,\theta)-\frac{1}{64}Z_{04}(r,\theta)$
$+Z_{11}(r,\theta)-\frac{1}{2}Z_{12}(r,\theta)+\frac{1}{4}Z_{13}(r,\theta)$
$-\frac{1}{8}Z_{14}(r,\theta)+Z_{21}(r,\theta)+Z_{31}(r,\theta)+Z_{41}(r,\theta))$
@@ 5 Discussion
In this work, we generalized periodic boundary conditions in Chen et al. (2024) to Dirichlet boundary conditions, demonstrating the adaptability of our approach to handle different types of constraints. This extension broadens the applicability of the method to a wider range of physical and mathematical problems, where Dirichlet boundary conditions play a critical role. The results underscore the potential of this generalized framework for accurately solving boundary value problems while maintaining computational efficiency and flexibility.
@@ 6 Acknowledgments
I would like to thank Professor Di Luo and Zhuo Chen, their valuable suggestions and great support throughout this project. This note was not reviewed, approved, or endorsed by Prof. Di Luo or Zhuo Chen. Any errors are solely the author’s responsibility. I would like to thank Xinjie He for her valuable contribution.